OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form . Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (wolfe8):

So the slope-intercept form is y=mx+b where m is the gradient/slope and b is the y-intercept. So what you need to do is rearrange the equation and then read from it the answers for 1.

OpenStudy (anonymous):

ok can you show me how i would do it plz

OpenStudy (wolfe8):

You have 2x + 3y = 1200 You want only y on the left side. So you will have to get rid of the 2x and scalar 3 on the left side. Do that, you will minus both sides with 2x to get 2x + 3y -2x = 1200 -2x 2y = 1200 - 2x Now all you need is to divide both sides with 2. Can you do that?

OpenStudy (anonymous):

yes 2y=600

OpenStudy (anonymous):

?

OpenStudy (agent0smith):

What happened to the -2x? It doesn't just disappear when you divide by 2. -2x/2 is...?

OpenStudy (wolfe8):

How familiar are you with algebra?

OpenStudy (anonymous):

i am familiar.. but not good

OpenStudy (anonymous):

-1

OpenStudy (wolfe8):

If I give you 2x+2y = 4 and tell you to divide it by two, what would you do exactly?

OpenStudy (anonymous):

i would divide 2x by 2 and 4 by 2

OpenStudy (wolfe8):

What about the 2y? When you divide an equation or expression by a number, you have to divide every single element in it by that number. You are basically doing \[(2x+2y=4)\times1/2\] Using the distributive rule, you will do 2x/2 + 2y/2 = 4/2 Same goes for multiplication. When doing addition and subtraction it goes as how I showed you in my workings earlier.

OpenStudy (wolfe8):

So can you try that again? 2x+2y = 4 divided by 2

OpenStudy (anonymous):

ok everything going to come down to x+2=2

OpenStudy (wolfe8):

Did you mistype that? Because when you divide a multiple of an unknown by a scalar, you can only do the process for the scalars. For example, as you did 2x/2 you only divide 2/2 and leave the x until the end of the process. So you get x. Dame goes for y, or z, or a or b or any unknown. Try again?

OpenStudy (agent0smith):

2y/2 = y. You can't make a variable just magically disappear.

OpenStudy (anonymous):

@agent0smith is this in slope intercept form?

OpenStudy (agent0smith):

is what...?